# Quarkiton: a one-quark state near a boundary of confinement phase of QCD

### v1

@@ Abstract
We discuss a one-quark state in the confinement phase near a reflective chromometallic boundary both at finite and zero temperature. Using numerical simulations of lattice Yang-Mills theory, we show that the test quark is confined to the neutral mirror by an attractive potential of the Cornell type, suggesting the existence of a mirror-bound one-quark state, a “quarkiton”. Surprisingly, the tension of the string spanned between the quark and the mirror is lower than the fundamental string tension. The quarkiton state exhibits a partial confinement: while the quark is localized in the vicinity of the mirror, it can still travel freely along it. Such quarkiton states share similarity with the surface excitons in metals and semiconductors that are bound to their negatively charged images at a boundary. The quarkitons can exist at the hadronic side of the phase interfaces in QCD that arise, for example, in the thermodynamic equilibrium of vortical quark–gluon plasma.
@@ I Introduction
A fundamental property of Quantum Chromodynamics (QCD) is color confinement—the empirical observation that isolated color-charged particles, such as individual quarks or gluons, are never detected under standard low-density, low-temperature environments in Nature Greensite (2020). Instead, all physically observable states are color-singlet combinations, manifesting themselves either as mesons (quark-antiquark pairs), baryons (three-quark states), or more exotic hadronic and gluonic (glueballs) configurations Bali and Schilling (1992). The confinement property disappears only under particular conditions, such as high temperatures or high baryonic densities Laermann and Philipsen (2003).
The confinement of color is a nonperturbative infrared phenomenon that cannot be revealed with the help of the standard perturbation theory in QCD. Confinement is believed to arise from the infrared dynamics of gluons, where the strong coupling constant becomes large, making perturbative methods inadequate. As a result, the Hilbert space of asymptotic states contains only color-neutral states. The confinement property is supported also by first-principle lattice simulations, which show that in the pure-glue non-Abelian gauge theories, the potential $V_{Q\bar{Q}}$ between a sufficiently separated static test quark $Q$ and anti-quark $\bar{Q}$ rises linearly Wilson (1974):
@@
The emergence of the linear potential is associated with the formation of a confining QCD string, which squeezes the electric color flux of quarks into a linear string-like structure with a finite energy density $\sigma$ per unit string length Greensite (2020).
Thus, the confinement property implies that an isolated, one-quark state is not a physical state. In the stringy picture of color confinement, this statement is very transparent. The quark state does not appear alone in the vacuum, as it always comes with a chromoelectric flux tube made of gluons. One end of the chromoelectric flux tube is attached to the source of the color field, the quark itself, while the other end goes essentially to infinity because there is no anti-quark for the flux tube to terminate at. Since the string carries certain energy per unit length, the total free energy of the system, “the quark plus the attached infinitely long string”, also becomes infinite, thus forbidding the existence of one-quark states in the confinement phase. Therefore, in the standard infinite-volume setting without boundaries, isolated one-quark states do not exist: the associated flux tube cannot terminate anywhere, its length becomes infinite and the total free energy associated with an isolated quark diverges.
@@
Figure 1: (a) A schematic visualization of the quarkiton state between a single quark $Q$ (right) and a neutral mirror (a vertical solid line). The quark is bound by a chromoelectric flux tube to the mirror by inducing an image antiquark $\bar{Q}$ (left). The quarkiton state (a) is a non-Abelian analog of the surface exciton (b) in which the electrons $(e^{-})$ and holes $(e^{+})$ form photon-mediated Coulomb-bound states with their images at the boundary.
Thus, a minimal energy requirement for a static quark located at a fixed distance from the chromoelectric mirror boundary implies that the chromoelectric flux tube, originating at the quark “QQ”, should terminate at the mirror “||” along the boundary normal. The string confines the quark to the mirror. Effectively, the heavy quark $Q$ is connected by the string to its negative image, an antiquark $\bar{Q}$. This configuration is visualized in Fig. 1(a)
Since the energy of the flux tube is linearly proportional to the length of the tube, this configuration would create a long-range confining linear potential between the quark and the mirror, separated by the distance $d$:
@@
The long-range attraction between the quark and the mirror is supplemented, at very short distances, by the strong repulsive potential of the mirror, which is impermeable to quarks. The quark should therefore form a single-quark state bound to the mirror, the “quarkiton” Chernodub et al. (2023a). A single-quark quarkiton state exhibits a partial color confinement since the quark is confined to the mirror by the flux tube and it cannot move far in the normal direction. However, the quarkiton can still freely propagate along the mirror in the transverse directions.
The quarkitons, along with their gluonic counterparts, “gluetons”, emerge naturally as boundary (edge) states in a non-Abelian version of the Casimir effect Chernodub et al. (2023a), which describes the interaction of neutral chromoelectric or chromomagnetic mirrors in a non-Abelian vacuum Chernodub et al. (2018); Karabali and Nair (2018); Chernodub et al. (2023a); Canfora et al. (2025); Dudal et al. (2025); Karabali et al. (2025).
@@
The potential between the electron $(e^{-})$ and its hole image $(e^{+})$ in the mirror is given by a Coulomb interaction, shown in Fig. 1(b). The electronic surface exciton is a bound electron–hole pair whose center-of-mass motion and wavefunction are confined to the vicinity of a surface or interface. As is clear from our description above, the quarkiton (glueton) excitation emerges essentially via the same physical mechanism: a quark (a gluon) coupled to its negative image in the mirror, an anti-quark (a gluon of a charge-conjugated color), and forms a neutral bound state. Another condensed-matter analogue of the non-Abelian bound states is given by fractional vortices in multi-component condensates, which are also linearly confined to the boundary Silaev (2011); Agterberg et al. (2014); Maiani et al. (2022).
@@
Let us stress that one of the constituents of the surface-exciton pair is a real particle (either an electron or a hole), while the other one (a hole or an electron, respectively) represents its image in the mirror. In QCD, analogously, the quarkiton is a single-quark state coupled to its anti-quark image that appears in the chromoelectric mirror (and likewise for a glueton). Both gluetons and quarkitons could potentially emerge as physical objects, as they can propagate along a reflective vacuum–hadronic-interior interface. Such interfaces arise naturally within an MIT bag model Chodos et al. (1974a, b). Also, stable confining-deconfining phase boundaries appear in rotating (quark-)gluon plasmas Chernodub (2021); Chernodub et al. (2023b); Braguta et al. (2024, 2025); Morales-Tejera et al. (2025).
@@
Certain properties of gluetons were studied in Ref. Chernodub et al. (2023a). It is instructive to compare the glueton mass with the mass of the glueball. The glueball is a color-neutral state made of gluons that exist in the bulk far from the boundaries. The mass of a ground-state gluon, by definition, defines a mass gap of the Yang-Mills system: there is no excitation with a mass lower than the ground-state gluon mass. The glueton is also a color-neutral boundary state, but it is formed by gluons bound to their image in a mirror so that the glueton is a boundary state with a restricted mobility. Unexpectedly, it appears that the mass of the glueton state is substantially lower than the mass of the glueball: the boundary glueton states exist below the mass gap set by the glueballs in the bulk Chernodub et al. (2023a).
@@
The “midgap” nature of the glueton state does not come as a surprise given that in solid-state physics, the states localized at the boundaries of a system (often called the edge states) can have lower masses than the bulk mass gap of the same system. This property is pertinent for the flat contacts of semiconductor structures (the Volkov-Pankratov states Volkov and Pankratov (1985)) as well as for the boundaries of topological insulators such as massless edge modes that appear in the spin Hall systems Hasan and Kane (2010); Qi and Zhang (2011). However, contrary to the mentioned boundary states, the glueton has a non-topological origin.
@@
Before going into the details, it is important to notice that the computation of the quark-mirror potential is not an easy or straightforward task at zero temperature since a single quark can only be introduced in the system with the help of the Polyakov loop $P$. This loop winds once around the imaginary time direction, which has an infinite length inversely proportional to a vanishing temperature $T\to 0$. Indeed, the free energy $F_{Q|}(T,d)$ of an isolated static color source located at a distance $d$ from the mirror is a finite quantity, so the expectation value of the Polyakov loop, $\langle P\rangle\sim e^{-F_{Q|}(T,d)/T}$ becomes exponentially suppressed Polyakov (1978). This severe suppression degrades the signal-to-noise ratio, complicating a reliable evaluation of the expectation value $\langle P\rangle$ at low temperatures. Therefore, in our work, we perform a calculation in a range of temperatures at the confinement phase and then make an extrapolation of the quark-mirror potential $F_{Q|}(T,R)$ to the zero-temperature limit, $T\to 0$. Preliminary results, at smaller lattices and with a different renormalization convention, were reported in Ref. Tanashkin et al. (2025).
The structure of this paper is as follows. In Sec. II, we discuss Yang-Mills theory on the lattice, show how to introduce the chromoelectric mirror, and describe the parameters of the numerical simulations. Section III is devoted to numerical calculation of the quark-mirror potential at finite temperature and to an extrapolation of the results to the zero-temperature case. In this section, we compare the quark-mirror potential with the quark-antiquark potential in the absence of the mirror in a broad range of temperatures. The last section summarizes our conclusions.
@@ II.1 Yang-Mills theory
In our paper, we use the standard Wilson formulation of the lattice version of the Yang-Mills action (3) which is suitable for numerical simulations:
@@
A static heavy quark is introduced at a spatial point $\bm{x}$} with the help of the Polyakov-loop operator:
@@ II.2 Chromometallic mirror
An ideal chromometallic mirror placed at the hypersurface $\mathcal{S}$} can be modeled by the gauge-invariant non-Abelian Casimir boundary conditions imposed on gluonic fields in SU($N$) gauge theory:
@@ II.3 Details of lattice simulations
We work on asymmetric lattices, $N_{t}\times N_{x}\times N_{y}\times N_{z}$. The physical temperature is inversely proportional to the temporal extension of the lattice, $N_{t}$ Gattringer and Lang (2010):
@@
The critical lattice coupling $\beta_{c}$ corresponding to the critical temperature of the deconfinement phase transition $T_{c}$ is a function of the lattice extension in the temporal direction $N_{t}$, cf. Eq. (7). The critical values of the coupling $\beta_{c}$ are listed, for various extensions $N_{t}$, in Table 1. Here, we used the results of Ref. Lucini et al. (2004) and also used our own numerical simulations to add more values that are needed for our detailed analysis.
@@
Table 1: Parameters of the numerical simulations. For each value of the temporal lattice extension $N_{s}$, we show the critical lattice coupling $\beta_{c}$ in the thermodynamic limit, the corresponding lattice spacing $a$ (in units of the zero-temperature string tension $\sigma_{0}$) and the range of the lattice spacings used (in physical units, fm). The last column shows the lowest temperature $T_{\rm min}$ (in units of the critical temperature $T_{c}$) used in our simulations for a particular value of $N_{t}$.
@@
The knowledge of the dependence of the lattice spacing $a(\beta)\sqrt{\sigma}_{0}$ from Ref. Athenodorou and Teper (2020) and the values of the critical values $\beta_{c}=\beta(N_{t})$ for each $N_{t}$, listed in Table 1, allows us to find the values of the coupling $\beta$ for other values of temperatures $T$ at all $N_{t}$. To this end, we use Eq. (7) and consider the ratio of $T_{c}/T$:
@@
We used spatially asymmetric lattices with $N_{x}=N_{y}>N_{z}$. We placed a flat static chromoelectric mirror at the spatial point $z=0$. It appears that the choice of the spatially transverse extension $N_{z}=30$ is enough to measure, with a satisfactory precision for our set of parameters, the effects of the mirror on the free energy of the quark. In order to achieve the thermodynamic limit, we studied the lattices with a set of large longitudinal extensions $N_{x}=N_{y}=132,192,224,254$. Together with the broad range of the lattice spacings, shown in Table 1, the variation of the spatial volume allowed us to achieve a continuum limit and ensure the independence of our results on ultraviolet (lattice spacing) and infrared (lattice size) parameters of the lattice simulations.
@@ III.1 Finite temperature results and renormalization
The expectation value of this operator is taken in a sufficiently large system, implying that the distance between the quark and the antiquark is much smaller than the size of the system in any spatial direction. In order to improve the statistics and enhance the signal-over-noise ratio, the expectation value (9) is averaged over all spatial points $\bm{x}$ and over all orientations of the inter-particle distance $\bm{R}$, compatible with the lattice symmetries.
@@
In our work, we introduce a static flat chromoelectric mirror at the position $z=0$ that extends through the whole system in the directions $x$ and $y$. The plate is, effectively, infinite, as it is closed via the periodic boundary conditions of a very large system. We use the lattice version of the perfect-chromometallic boundary condition (6) following the procedure given in Section II.2.
@@
Figure 2: Unrenormalized free energy of a static heavy quark as a function of distance to the mirror $d$ at temperature $T=0.629T_{c}$ for several values of the temporal extension $N_{t}$ of the lattice, which correspond, via Eq. (7), to different ultraviolet cutoffs (lattice spacings) $a$. The linear slopes of potentials are marked by the straight lines. The physical scale is given by the zero-temperature fundamental string tension $\sigma_{0}$. The calculation is performed at $N_{t}\times 254\times 254\times 30$ lattices.
In Fig. 2, we show an example of the single-quark free energy near the mirror, which is computed numerically using the single-point Polyakov loop (10) at the temperature $T=0.629T_{c}$. We present the data for various extensions of the lattice in the temporal direction $N_{t}$, which allows us to calculate the same physical quantity at different lattice spacings $a$. Our numerical data shows that while the quark-mirror potential is affected by the ultraviolet cutoff $a$, the linear coefficients of the straight slopes, marked by the straight lines in Fig. 2, are largely independent of the value of the lattice spacing $a$. Also, at a sufficiently large separation $d$ from the mirror, the single-quark free energy flattens and collapses to the same value for all values of the cutoff $a$. The value of the asymptotic Polyakov loop depends on the spatial size of the system. The $d\to\infty$ plateau vanishes in the thermodynamic limit in the consistency of the color-confining property.
Figure 3: The renormalized free energy of a single heavy quark at the distance $d$ to the chromoelectric mirror at the temperatures (a) $T=0.629T_{c}$ and (b) $T=0.996T_{c}$ for various lattice spacings controlled by the temporal extension of the lattice $N_{t}$. The solid lines show the best fits by the Cornell potential (11). We use the $N_{t}\times 254\times 254\times 30$ lattices.
Figure 2 indicates that the single-quark potential is affected by the ultraviolet effects that provide an additive correction to the quark self-energy. Therefore, the free energy needs to be renormalized similarly to the renormalization of the quark-antiquark potential in the absence of the boundary Kaczmarek et al. (2004). In our paper, we perform the renormalization of the quark-mirror interaction following the idea of matching the low-distance behavior of the potentials discussed in Ref. Kaczmarek et al. (2004).
@@
After subtracting the ultraviolet contribution from the lattice data and excluding the flattening long-distance tails, we finally get the renormalized free energy $F_{Q|}^{\mathrm{ren}}(d)$. In Fig. 3 we show the examples of the renormalized free energies of quarks near the mirror for two representative temperatures, far from the transition point, $T=0.629T_{c}$ and very close to the transition, $T=0.996T_{c}$.
The data for the renormalized free energy perfectly arranges to the single curve, thus indicating the independence of our data on the ultraviolet lattice cutoff $a=a(\beta)$. We also checked the insensitivity of the results on the spatial lattice size $L_{s}=N_{s}a$.
@@
The second term in Eq. (11) corresponds to the non-perturbative linear ($\propto d$) growth of the free energy due to the formation of the confining string between the heavy quark and the neutral mirror. This term is important at large distances. For the data shown in Fig. 3, the best-fit values for the string tension in Eq. (11) are $\sigma_{Q|}=0.688(10)\sigma_{0}$ for $T=0.629T_{c}$ and $\sigma_{Q|}=0.136(8)\sigma_{0}$ for $T=0.996T_{c}$. The last term in Eq. (11) is an inessential additive factor.
The excellent fits of the free energy of an isolated quark near a neutral mirror by the Cornell potential (11) indicate that the quark indeed generates an anti-quark image in a mirror and gets attracted to it. At large distances from the mirror $d$, the linear attraction of the heavy test quark to the mirror implies that the quark is indeed confined to the chromoelectric mirror, thus supporting our previous conclusion on the formation of the non-perturbative quarkiton boundary state.

### v2

# One-quark state near a boundary of the confinement phase of QCD

@@ Abstract
We discuss a one-quark state in the confinement phase near a reflective chromometallic boundary both at finite and zero temperature. Using numerical simulations of lattice Yang-Mills theory, we show that the test quark is confined to the neutral mirror by an attractive potential of the Cornell type, suggesting the existence of a mirror-bound one-quark state, a “quarkiton.” Surprisingly, the tension of the string spanned between the quark and the mirror is lower than the fundamental string tension. The quarkiton state exhibits a partial confinement: while the quark is localized in the vicinity of the mirror, it can still travel freely along it. Such quarkiton states share similarity with the surface excitons in metals and semiconductors that are bound to their negatively charged images at a boundary. The quarkitons can exist at the hadronic side of the phase interfaces in QCD that arise, for example, in the thermodynamic equilibrium of vortical quark–gluon plasma.
@@ I Introduction
A fundamental property of QCD is color confinement—the empirical observation that isolated color-charged particles, such as individual quarks or gluons, are never detected under standard low-density, low-temperature environments in Nature Greensite (2020). Instead, all physically observable states are color-singlet combinations, manifesting themselves either as mesons (quark-antiquark pairs), baryons (three-quark states), or more exotic hadronic and gluonic (glueballs) configurations Bali and Schilling (1992). The confinement property disappears only under particular conditions, such as high temperatures or high baryonic densities Laermann and Philipsen (2003).
The confinement of color is a nonperturbative IR phenomenon that cannot be revealed with the help of the standard perturbation theory in QCD. Confinement is believed to arise from the IR dynamics of gluons, where the strong coupling constant becomes large, making perturbative methods inadequate. As a result, the Hilbert space of asymptotic states contains only color-neutral states. The confinement property is supported also by first-principle lattice simulations, which show that in the pure-glue non-Abelian gauge theories, the potential $V_{Q\bar{Q}}$ between a sufficiently separated static test quark $Q$ and antiquark $\bar{Q}$ rises linearly Wilson (1974), as follows:
@@
The emergence of the linear potential is associated with the formation of a confining QCD string, which squeezes the electric color flux of quarks into a linear stringlike structure with a finite energy density $\sigma$ per unit string length Greensite (2020).
Thus, the confinement property implies that an isolated, one-quark state is not a physical state. In the stringy picture of color confinement, this statement is very transparent. The quark state does not appear alone in the vacuum, as it always comes with a chromoelectric flux tube made of gluons. One end of the chromoelectric flux tube is attached to the source of the color field, the quark itself, while the other end goes essentially to infinity because there is no antiquark at which the flux tube can terminate. Since the string carries certain energy per unit length, the total free energy of the system, “the quark plus the attached infinitely long string,” also becomes infinite, thus forbidding the existence of one-quark states in the confinement phase. Therefore, in the standard infinite-volume setting without boundaries, isolated one-quark states do not exist: the associated flux tube cannot terminate anywhere, its length becomes infinite, and the total free energy associated with an isolated quark diverges.
@@
Figure 1: (a) Schematic visualization of the quarkiton state between a single quark $Q$ (right) and a neutral mirror (a vertical solid line). The quark is bound by a chromoelectric flux tube to the mirror by inducing an image antiquark $\bar{Q}$ (left). The quarkiton state (a) is a non-Abelian analog of the surface exciton (b) in which the electrons $(e^{-})$ and holes $(e^{+})$ form photon-mediated Coulomb-bound states with their images at the boundary.
Thus, a minimal energy requirement for a static quark located at a fixed distance from the chromoelectric mirror boundary implies that the chromoelectric flux tube, originating at the quark “QQ,” should terminate at the mirror “||” along the boundary normal. The string confines the quark to the mirror. Effectively, the heavy quark $Q$ is connected by the string to its negative image, an antiquark $\bar{Q}$. This configuration is visualized in Fig. 1(a)
Since the energy of the flux tube is linearly proportional to the length of the tube, this configuration would create a long-range confining linear potential between the quark and the mirror, separated by the distance $d$, as follows:
@@
The long-range attraction between the quark and the mirror is supplemented, at very short distances, by the strong repulsive potential of the mirror, which is impermeable to quarks. The quark should therefore form a single-quark state bound to the mirror, the quarkiton Chernodub et al. (2023a). A single-quark quarkiton state exhibits a partial color confinement since the quark is confined to the mirror by the flux tube and it cannot move far in the normal direction. However, the quarkiton can still freely propagate along the mirror in the transverse directions.
The quarkitons, along with their gluonic counterparts, “gluetons,” emerge naturally as boundary (edge) states in a non-Abelian version of the Casimir effect Chernodub et al. (2023a), which describes the interaction of neutral chromoelectric or chromomagnetic mirrors in a non-Abelian vacuum Chernodub et al. (2018); Karabali and Nair (2018); Chernodub et al. (2023a); Canfora et al. (2025); Dudal et al. (2025); Karabali et al. (2025).
@@
The potential between the electron $(e^{-})$ and its hole image $(e^{+})$ in the mirror is given by a Coulomb interaction, shown in Fig. 1(b). The electronic surface exciton is a bound electron–hole pair whose c.m. motion and wave function are confined to the vicinity of a surface or interface. As is clear from our description above, the quarkiton (glueton) excitation emerges essentially via the same physical mechanism: a quark (a gluon) couples to its negative image in the mirror, an antiquark (a gluon of a charge-conjugated color), and forms a neutral bound state. Another condensed-matter analog of the non-Abelian bound states is given by fractional vortices in multicomponent condensates, which are also linearly confined to the boundary Silaev (2011); Agterberg et al. (2014); Maiani et al. (2022).
@@
Let us stress that one of the constituents of the surface-exciton pair is a real particle (either an electron or a hole), while the other one (a hole or an electron, respectively) represents its image in the mirror. In QCD, analogously, the quarkiton is a single-quark state coupled to its antiquark image that appears in the chromoelectric mirror (and likewise for a glueton). Both gluetons and quarkitons could potentially emerge as physical objects, as they can propagate along a reflective vacuum–hadronic-interior interface. Such interfaces arise naturally within an MIT bag model Chodos et al. (1974a, b). Also, stable confining-deconfining phase boundaries appear in rotating (quark-)gluon plasmas Chernodub (2021); Chernodub et al. (2023b); Braguta et al. (2024, 2025); Morales-Tejera et al. (2025).
@@
Certain properties of gluetons were studied in Ref. Chernodub et al. (2023a). It is instructive to compare the glueton mass with the mass of the glueball. The glueball is a color-neutral state made of gluons that exist in the bulk far from the boundaries. The mass of a ground-state glueball, by definition, defines a mass gap of the Yang-Mills system: there is no excitation with a mass lower than the ground-state glueball mass. The glueton is also a color-neutral boundary state, but it is formed by gluons bound to their image in a mirror so that the glueton is a boundary state with a restricted mobility. Unexpectedly, it appears that the mass of the glueton state is substantially lower than the mass of the glueball: the boundary glueton states exist below the mass gap set by the glueballs in the bulk Chernodub et al. (2023a).
@@
The “midgap” nature of the glueton state does not come as a surprise, given that in solid-state physics, the states localized at the boundaries of a system (often called the edge states) can have lower masses than the bulk mass gap of the same system. This property is pertinent for the flat contacts of semiconductor structures (the Volkov-Pankratov states Volkov and Pankratov (1985)) as well as for the boundaries of topological insulators such as massless edge modes that appear in the spin Hall systems Hasan and Kane (2010); Qi and Zhang (2011). However, contrary to the mentioned boundary states, the glueton has a nontopological origin.
@@
Before going into the details, it is important to notice that the computation of the quark-mirror potential is not an easy or straightforward task at zero temperature, since a single quark can only be introduced in the system with the help of the Polyakov loop $P$. This loop winds once around the imaginary time direction, which has an infinite length inversely proportional to a vanishing temperature $T\to 0$. Indeed, the free energy $F_{Q|}(T,d)$ of an isolated static color source located at a distance $d$ from the mirror is a finite quantity, so the expectation value of the Polyakov loop, $\langle P\rangle\sim e^{-F_{Q|}(T,d)/T}$, becomes exponentially suppressed Polyakov (1978). This severe suppression degrades the SNR, complicating a reliable evaluation of the expectation value $\langle P\rangle$ at low temperatures. Therefore, in our work, we perform a calculation in a range of temperatures at the confinement phase and then make an extrapolation of the quark-mirror potential $F_{Q|}(T,R)$ to the zero-temperature limit, $T\to 0$. Preliminary results, at smaller lattices and with a different renormalization convention were reported in Ref. Tanashkin et al. (2025).
The structure of this paper is as follows: in Sec. II, we discuss Yang-Mills theory on the lattice, show how to introduce the chromoelectric mirror, and describe the parameters of the numerical simulations. Section III is devoted to numerical calculation of the quark-mirror potential at finite temperature and to an extrapolation of the results to the zero-temperature case. In this section, we compare the quark-mirror potential with the quark-antiquark potential in the absence of the mirror in a broad range of temperatures. The last section summarizes our conclusions.
@@ II.1 Yang-Mills theory
In our paper, we use the standard Wilson formulation of the lattice version of the Yang-Mills action (3) which is suitable for numerical simulations, as follows:
@@
A static heavy quark is introduced at a spatial point $\bm{x}$} with the help of the Polyakov-loop operator,
@@ II.2 Chromometallic mirror
An ideal chromometallic mirror placed at the hypersurface $\mathcal{S}$} can be modeled by the gauge-invariant non-Abelian Casimir boundary conditions imposed on gluonic fields in SU($N$) gauge theory, as follows:
@@ II.3 Details of lattice simulations
We work on asymmetric lattices, $N_{t}\times N_{x}\times N_{y}\times N_{z}$. The physical temperature is inversely proportional to the temporal extension of the lattice, $N_{t}$ Gattringer and Lang (2010), as follows:
@@
The critical lattice coupling $\beta_{c}$, corresponding to the critical temperature of the deconfinement phase transition $T_{c}$ is a function of the lattice extension in the temporal direction $N_{t}$, cf., Eq. (7). The critical values of the coupling $\beta_{c}$ are listed, for various extensions $N_{t}$, in Table 1. Here, we used the results of Ref. Lucini et al. (2004) and also used our own numerical simulations to add more values that are needed for our detailed analysis.
@@
Table 1: Parameters of the numerical simulations. For each value of the temporal lattice extension $N_{s}$, we show the critical lattice coupling $\beta_{c}$ in the thermodynamic limit, the corresponding lattice spacing $a$ (in units of the zero-temperature string tension $\sigma_{0}$), and the range of the lattice spacings used (in physical units, fm). The last column shows the lowest temperature $T_{\rm min}$ (in units of the critical temperature $T_{c}$) used in our simulations for a particular value of $N_{t}$.
@@
The knowledge of the dependence of the lattice spacing $a(\beta)\sqrt{\sigma}_{0}$ from Ref. Athenodorou and Teper (2020) and the values of the critical values $\beta_{c}=\beta(N_{t})$ for each $N_{t}$, listed in Table 1, allow us to find the values of the coupling $\beta$ for other values of temperatures $T$ at all $N_{t}$. To this end, we use Eq. (7) and consider the ratio of $T_{c}/T$, as follows:
@@
We used spatially asymmetric lattices with $N_{x}=N_{y}>N_{z}$. We placed a flat static chromoelectric mirror at the spatial point $z=0$. It appears that the choice of the spatially transverse extension $N_{z}=30$ is enough to measure, with a satisfactory precision for our set of parameters, the effects of the mirror on the free energy of the quark. In order to achieve the thermodynamic limit, we studied the lattices with a set of large longitudinal extensions $N_{x}=N_{y}=132,192,224,254$. Together with the broad range of the lattice spacings, shown in Table 1, the variation of the spatial volume allowed us to achieve a continuum limit and ensure the independence of our results on UV (lattice spacing) and IR (lattice size) parameters of the lattice simulations.
@@ III.1 Finite temperature results and renormalization
The expectation value of this operator is taken in a sufficiently large system, implying that the distance between the quark and the antiquark is much smaller than the size of the system in any spatial direction. In order to improve the statistics and enhance the SNR ratio, the expectation value (9) is averaged over all spatial points $\bm{x}$ and over all orientations of the interparticle distance $\bm{R}$, compatible with the lattice symmetries.
@@
In our work, we introduce a static flat chromoelectric mirror at the position $z=0$ that extends through the whole system in the directions $x$ and $y$. The plate is, effectively, infinite, as it is closed via the periodic boundary conditions of a very large system. We use the lattice version of the perfect-chromometallic boundary condition (6) following the procedure given in Sec. II.2.
@@
Figure 2: Unrenormalized free energy of a static heavy quark as a function of distance to the mirror $d$ at temperature $T=0.629T_{c}$ for several values of the temporal extension $N_{t}$ of the lattice, which correspond, via Eq. (7), to different UV cutoffs (lattice spacings) $a$. The linear slopes of potentials are marked by the straight lines. The physical scale is given by the zero-temperature fundamental string tension $\sigma_{0}$. The calculation is performed at $N_{t}\times 254\times 254\times 30$ lattices.
In Fig. 2, we show an example of the single-quark free energy near the mirror, which is computed numerically using the single-point Polyakov loop (10) at the temperature $T=0.629T_{c}$. We present the data for various extensions of the lattice in the temporal direction $N_{t}$, which allows us to calculate the same physical quantity at different lattice spacings $a$. Our numerical data show that while the quark-mirror potential is affected by the UV cutoff $a$, the linear coefficients of the straight slopes, marked by the straight lines in Fig. 2, are largely independent of the value of the lattice spacing $a$. Also, at a sufficiently large separation $d$ from the mirror, the single-quark free energy flattens and collapses to the same value for all values of the cutoff $a$. The value of the asymptotic Polyakov loop depends on the spatial size of the system. The $d\to\infty$ plateau vanishes in the thermodynamic limit in the consistency of the color-confining property.
Figure 3: Renormalized free energy of a single heavy quark at the distance $d$ to the chromoelectric mirror at the temperatures (a) $T=0.629T_{c}$ and (b) $T=0.996T_{c}$ for various lattice spacings controlled by the temporal extension of the lattice $N_{t}$. The solid lines show the best fits by the Cornell potential (11). We use the $N_{t}\times 254\times 254\times 30$ lattices.
Figure 2 indicates that the single-quark potential is affected by the UV effects that provide an additive correction to the quark self-energy. Therefore, the free energy needs to be renormalized similarly to the renormalization of the quark-antiquark potential in the absence of the boundary Kaczmarek et al. (2004). In our paper, we perform the renormalization of the quark-mirror interaction following the idea of matching the low-distance behavior of the potentials discussed in Ref. Kaczmarek et al. (2004).
@@
After subtracting the UV contribution from the lattice data and excluding the flattening long-distance tails, we finally get the renormalized free energy $F_{Q|}^{\mathrm{ren}}(d)$. In Fig. 3, we show the examples of the renormalized free energies of quarks near the mirror for two representative temperatures, far from the transition point, $T=0.629T_{c}$ and very close to the transition, $T=0.996T_{c}$.
The data for the renormalized free energy perfectly arrange to the single curve, thus indicating the independence of our data on the UV lattice cutoff $a=a(\beta)$. We also checked the insensitivity of the results on the spatial lattice size $L_{s}=N_{s}a$.
@@
The second term in Eq. (11) corresponds to the nonperturbative linear ($\propto d$) growth of the free energy due to the formation of the confining string between the heavy quark and the neutral mirror. This term is important at large distances. For the data shown in Fig. 3, the best-fit values for the string tension in Eq. (11) are $\sigma_{Q|}=0.688(10)\sigma_{0}$ for $T=0.629T_{c}$ and $\sigma_{Q|}=0.136(8)\sigma_{0}$ for $T=0.996T_{c}$. The last term in Eq. (11) is an inessential additive factor.
The excellent fits of the free energy of an isolated quark near a neutral mirror by the Cornell potential (11) indicate that the quark indeed generates an antiquark image in a mirror and gets attracted to it. At large distances from the mirror $d$, the linear attraction of the heavy test quark to the mirror implies that the quark is indeed confined to the chromoelectric mirror, thus supporting our previous conclusion on the formation of the nonperturbative quarkiton boundary state.